# A Survey on LLM-as-a-Judge

### v1

@@ 2.3.1. Extracting specific tokens
As showed in In-context Learning (Section 2.1), when the evaluation target take the form of a score, selecting specific options, or responding with Yes/No, applying rule-match to extract the corresponding token from the response generated during probability distribution iteration is common used. It is worth noting that Yes/No is a broad definition, including custom statements involving judgment. Considering a Yes/No question for evaluation in custom phrases (Tian et al., 2023b): "Modification needed." and "No modification needed." or a yes-no question "Does the above answer need to be further modified?". When the input sample is put through the template, it might have outputs such as ”Modification needed.”, ”Conclusion: Modification needed.” or ”Yes”. This variance in response formats is difficult to parse consistently. The corresponding post-processing with the response is necessary. Using rules to extract specific tokens for our designed prompts and input content, as well as the backbone model used for the evaluator, all have higher requirements as we discussed in Section 2.2. In contextual learning, if there is no clear indication of the output format for response, there may be various expressions of evaluation, which can be seen in Figure 2. For example, ”Response 1 is better” and ”The better one is response 1”, which convey the same choice but differ in format leading to the difficulty of rule recognition. Simple solutions often involve providing clear instructions, such as ”The last sentence should be started with ’The better response is’”, or using a few-shot strategy. Also, the general model with insufficient instruction following capability may not be able to generate the evaluation format and content of the target according to the instruction, resulting in the post-processing extracted according to the rules not as smooth as expected.
@@ 2.4. Evaluation Pipeline
There are three common scenarios for using LLM-as-a-Judge evaluation pipelines showed in Figure 2, which are LLM-as-a-Judge for LLMs, LLM-as-a-Judge for data, and LLM-as-a-Judge for agent respectively.
@@ 2.4.1. LLM-as-a-Judge for model
Recent advancements in using Large Multimodal Models (LMMs) as evaluators have showcased their potential to perform complex judgment tasks in vision-language scenarios. Proprietary models like GPT-4V and GPT-4o have been pivotal in benchmarks such as detailed captioning and visual chats, utilizing both pointwise and pairwise evaluation methods (Liu et al., 2024a; Zhang et al., 2023b; Lu et al., 2024). Open-source alternatives have emerged, with Prometheus-Vision (Lee et al., 2024a) being the first vision-language model specifically trained to act as an evaluator for user-designed scoring criteria. While Prometheus-Vision introduced the concept of open-source evaluators with a focus on specialized tasks, it remains limited to predefined criteria. In contrast, LLaVA-Critic (Xiong et al., 2024), another open-source innovation, expands the scope by serving as a generalist evaluator. Trained on diverse and detailed datasets, LLaVA-Critic provides robust scoring and preference learning, closely aligning with human and proprietary evaluations. These models mark significant progress in democratizing and enhancing multimodal evaluation tools.
@@ 2.4.2. LLM-as-a-Judge for data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022a; Ramamurthy et al., 2022). The most successful way is applying a RLHF framework (Ouyang et al., 2022a) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023c) and principle-driven self-alignment (Sun et al., 2023b).
@@
Recent research on evaluating multimodal data focuses on addressing vision-language misalignments in Multimodal Large Language Models (MLLMs), which often cause hallucinations—outputs inconsistent with visual or contextual evidence (Li et al., 2023a; Wang et al., 2023e; Cui et al., 2023). Techniques like Reinforcement Learning from Human Feedback (RLHF) and Factually Augmented RLHF have been employed to improve model alignment by incorporating structured ground-truth data and image captions, enhancing hallucination detection (Sun et al., 2023a). Benchmarks such as MLLM-as-a-Judge (Chen et al., 2024b) assess these models using tasks like scoring, pair comparison, and batch ranking, revealing limitations in alignment with human preferences. Persistent issues include biases (e.g., position, verbosity) and hallucinations, with even advanced models like GPT-4V displaying challenges. While pair comparison tasks align better with human judgment, scoring and batch ranking require significant improvements for reliable deployment. These findings emphasize the need for innovative frameworks and datasets to refine MLLM evaluation and alignment.
@@ 2.4.3. LLM-as-a-Judge for agent
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023; Shinn et al., 2024). Both approaches are briefly illustrated in Figure 7. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2024) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.
@@ 3. Improvement Strategy
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024b) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c) / SocREval (He et al., 2024b) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024b) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024a) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024c)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c) / LLM-EVAL (Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
Evaluation Datasets, text width=7em [PandaLM (Wang et al., 2023d) / SALAD-Bench (Li et al., 2024b) / OffsetBias (Park et al., 2024) / JudgeLM (Zhu et al., 2023a) / CritiqueLLM (Ke et al., 2024)] ] [Iterative Optimization
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023d) / Bai et al. (Bai et al., 2024b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1. Design Strategy of Evaluation Prompts
An evaluation prompt is an input to LLM evaluators, which is used to guide the LLMs to complete the required evaluation tasks. LLMs possess in-context learning ability to learn how to perform specified tasks through relevant examples or instructions provided in prompts without requiring weight updates or retraining(Brown, 2020). It indicates that the design strategy of evaluation prompts will significantly impact the effectiveness of LLM-as-a-judge. Therefore, how to optimize the design of evaluation prompts, including better methods to help LLMs understand the evaluation tasks and produce evaluation results, is the most direct and effective way to improve the evaluation performance of LLM-as-a-judge.
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown, 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024b) and GPTScore(Fu et al., 2024).
@@
In addition to providing hight-quality examples for LLMs to inference, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly including the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024c) use Chain-of-Thought(CoT)(Wei et al., 2022) to provide guidance for LLMs. SocREval(He et al., 2024b) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. proposes Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024b) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition and thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024a) analyzed and validated the impact of position bias on LLM-as-a-judge, and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@
To address the challenge of LLMs’ absolute scoring being less robust than relative comparing(Raina et al., 2024), some research works convert scoring tasks into pairwise comparison, thereby enhancing the reliability of evaluation results. Liu et al.(Liu et al., 2024c) transform the scoring evaluation to ranking evaluation and introduce Pairwise-Preference Search (PARIS), which employs LLMs to conduct pairwise comparisons locally and efficiently ranks candidate texts globally, making evaluation results more aligned with human preferences.

### v2

@@ 2.3. Post-processing Method
We will provide a detailed explanation of these methods. However, it is important to note that each method has significant limitations when evaluating objective questions. For example, in text response evaluation (Yu et al., 2024b), failing to accurately extract the key answer token from the LLM’s response can result in incorrect evaluation outcomes. These challenges in post-processing are deeply connected to the prompt design used in the earlier ICL stages and the selected model’s ability to follow instructions reliably.
@@ 2.3.1. Extracting specific tokens
As showed in In-context Learning (Section 2.1), when the evaluation target take the form of a score, selecting specific options, or responding with Yes/No, applying rule-match to extract the corresponding token from the response generated during probability distribution iteration is common used. It is worth noting that Yes/No is a broad definition, including custom statements involving judgment. Considering a Yes/No question for evaluation in custom phrases (Tian et al., 2024): "Modification needed." and "No modification needed." or a yes-no question "Does the above answer need to be further modified?". When the input sample is put through the template, it might have outputs such as ”Modification needed.”, ”Conclusion: Modification needed.” or ”Yes”. This variance in response formats is difficult to parse consistently. The corresponding post-processing with the response is necessary. Using rules to extract specific tokens for our designed prompts and input content, as well as the backbone model used for the evaluator, all have higher requirements as we discussed in Section 2.2. In contextual learning, if there is no clear indication of the output format for response, there may be various expressions of evaluation, which can be seen in Figure 2. For example, ”Response 1 is better” and ”The better one is response 1”, which convey the same choice but differ in format leading to the difficulty of rule recognition. Simple solutions often involve providing clear instructions, such as ”The last sentence should be started with ’The better response is’”, or using a few-shot strategy. Also, the general model with insufficient instruction following capability may not be able to generate the evaluation format and content of the target according to the instruction, resulting in the post-processing extracted according to the rules not as smooth as expected.
@@ 2.4. Evaluation Pipeline
After completing the three processes, we obtain the final evaluation. From input to output, these steps collectively form the LLM-as-a-Judge evaluation pipeline, as illustrated in Figure 2. This pipeline is commonly applied in three scenarios: LLM-as-a-Judge for LLMs, LLM-as-a-Judge for data, and LLM-as-a-Judge for agents.
@@ 2.4.1. LLM-as-a-Judge for model
Recent advancements in using Large Multimodal Models (LMMs) as evaluators have showcased their potential to perform complex judgment tasks in vision-language scenarios. Proprietary models like GPT-4V and GPT-4o have been pivotal in benchmarks such as detailed captioning and visual chats, utilizing both pointwise and pairwise evaluation methods (Liu et al., 2023b; Zhang et al., 2023b; Lu et al., 2024). Open-source alternatives have emerged, with Prometheus-Vision (Lee et al., 2024a) being the first vision-language model specifically trained to act as an evaluator for user-designed scoring criteria. While Prometheus-Vision introduced the concept of open-source evaluators with a focus on specialized tasks, it remains limited to predefined criteria. In contrast, LLaVA-Critic (Xiong et al., 2024), another open-source innovation, expands the scope by serving as a generalist evaluator. Trained on diverse and detailed datasets, LLaVA-Critic provides robust scoring and preference learning, closely aligning with human and proprietary evaluations. These models mark significant progress in democratizing and enhancing multimodal evaluation tools.
@@ 2.4.2. LLM-as-a-Judge for data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022c; Ramamurthy et al., 2023). The most successful way is applying a RLHF framework (Ouyang et al., 2022c) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023d) and principle-driven self-alignment (Sun et al., 2023c).
@@
However, solely relying on LLM-as-a-Judge for data annotation poses challenges, particularly as the value of annotated data diminishes with the rapid improvement of model performance. To address this, approaches like Self-Taught Evaluator (Wang et al., 2024a) offer a promising alternative by eliminating the need for human annotations. This method leverages synthetic training data, starting with unlabeled instructions and generating contrasting outputs from models. These outputs are then used to train an LLM-as-a-Judge to produce reasoning traces and final judgments. With each iteration, the evaluator improves by learning from its refined predictions, creating a cycle of continuous self-enhancement. This iterative approach not only keeps annotations relevant but also ensures that evaluators evolve alongside advancing models.
Recent research on evaluating multimodal data focuses on addressing vision-language misalignments in Multimodal Large Language Models (MLLMs), which often cause hallucinations—outputs inconsistent with visual or contextual evidence (Li et al., 2023a; Wang et al., 2023e; Cui et al., 2023). Techniques like Reinforcement Learning from Human Feedback (RLHF) and Factually Augmented RLHF have been employed to improve model alignment by incorporating structured ground-truth data and image captions, enhancing hallucination detection (Sun et al., 2023b). Benchmarks such as MLLM-as-a-Judge (Chen et al., 2024b) assess these models using tasks like scoring, pair comparison, and batch ranking, revealing limitations in alignment with human preferences. Persistent issues include biases (e.g., position, verbosity) and hallucinations, with even advanced models like GPT-4V displaying challenges. While pair comparison tasks align better with human judgment, scoring and batch ranking require significant improvements for reliable deployment. These findings emphasize the need for innovative frameworks and datasets to refine MLLM evaluation and alignment.
@@ 2.4.3. LLM-as-a-Judge for agent
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023; Shinn et al., 2023). Both approaches are briefly illustrated in Figure 7. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2023) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.
@@ 3. Improvement Strategy
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024c) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / SocREval (He et al., 2024) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024a) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024b) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024b)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
Evaluation Datasets, text width=7em [PandaLM (Wang et al., 2023d) / SALAD-Bench (Li et al., 2024c) / OffsetBias (Park et al., 2024) / JudgeLM (Zhu et al., 2023a) / CritiqueLLM (Ke et al., 2024)] ] [Iterative Optimization
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023e) / Bai et al. (Bai et al., 2023b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1. Design Strategy of Evaluation Prompts
An evaluation prompt is an input to LLM evaluators, which is used to guide the LLMs to complete the required evaluation tasks. LLMs possess in-context learning ability to learn how to perform specified tasks through relevant examples or instructions provided in prompts without requiring weight updates or retraining(Brown et al., 2020). It indicates that the design strategy of evaluation prompts will significantly impact the effectiveness of LLM-as-a-judge. Therefore, how to optimize the design of evaluation prompts, including better methods to help LLMs understand the evaluation tasks and produce evaluation results, is the most direct and effective way to improve the evaluation performance of LLM-as-a-judge.
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown et al., 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024c) and GPTScore(Fu et al., 2024).
@@
In addition to providing hight-quality examples for LLMs to inference, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly including the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024d) use Chain-of-Thought(CoT)(Wei et al., 2022) to provide guidance for LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. proposes Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition and thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024b) analyzed and validated the impact of position bias on LLM-as-a-judge, and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@
To address the challenge of LLMs’ absolute scoring being less robust than relative comparing(Raina et al., 2024), some research works convert scoring tasks into pairwise comparison, thereby enhancing the reliability of evaluation results. Liu et al.(Liu et al., 2024b) transform the scoring evaluation to ranking evaluation and introduce Pairwise-Preference Search (PARIS), which employs LLMs to conduct pairwise comparisons locally and efficiently ranks candidate texts globally, making evaluation results more aligned with human preferences.